OpenStudy (anonymous):

Find the solution set (2x-1)^2-3(2x-1)=-2

OpenStudy (anonymous):

\[(2x-1)^2-3(2x-1)^2+2=0\] solve \[w^2-3w+2=0\] then replace \(w\) by \(2x-1\) and solve for \(x\)

OpenStudy (anonymous):

pretty easy in this case because \[w^2-3w+2=0\] factors as \[(w-2)(w-1)=0\] making \[w=1\] or \[w=2\] so your only job now is to solve \[2x-1=1\] and \[2x-1=2\] for \(x\)

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

you good from there right?

OpenStudy (anonymous):

yes, thank you :) ill try to solve it on my own

OpenStudy (anonymous):

x=1 and x=3/2